OpenStudy (anonymous):

will medal and fan solve for x x-3 -----=5 2 a.–4 b.–7 c.–16 d.–13

OpenStudy (anonymous):

Are you asking:\[\frac{ x - 3 }{ 2 } = 5\]

OpenStudy (anonymous):

yes

OpenStudy (legends):

Ok so first we find 5*2

OpenStudy (anonymous):

10

OpenStudy (shadeyghost):

then add 3

OpenStudy (legends):

ok so x-3=10

OpenStudy (anonymous):

@Legends Try to explain it a bit?

OpenStudy (anonymous):

13

OpenStudy (shadeyghost):

there is your answer

OpenStudy (anonymous):

so -13

OpenStudy (shadeyghost):

wait, that doesn't make sense, how is it a negative number?

OpenStudy (legends):

I know thats what im thinkning

OpenStudy (anonymous):

\(\dfrac{x-3}{2} = 5\) First we multiply 2 to both sides.. 5 * 2 = 10 you said.. \(x - 3 = 10\) Then we add 3 to both sides, 10 + 3 is 13 you said.. \(x = 13\)

OpenStudy (anonymous):

The answer is NOT negative 13. 13 is positive. @briggs

OpenStudy (shadeyghost):

but that isn't a choice, all of the answer choices are negative

OpenStudy (anonymous):

He probably made a mistake in his equation.

OpenStudy (legends):

Yea but it says -13 i think you got it wrong @briggs

OpenStudy (anonymous):

@briggs You probably made a mistake in the equation.

OpenStudy (anonymous):

@King.Void. I can't message you back..

OpenStudy (anonymous):

that is what o the practice unless he teachers got it wrong

OpenStudy (anonymous):

I would ask your teacher, there's no way it's -13.